### v1

# Photo-nuclear cross sections on Au

Abstract
A method was developed for measuring photonuclear reactions concurrently at several discrete photon beam energies on a stack of different target materials via a single irradiation. Concentric ring targets of the materials (in order from front to back targets: Au, TiO, Zn, Os, and Au) were irradiated at the High Intensity Gamma-ray Source (HIS). As a proof of principle, we report the result of the cross section measurements from the front Au target. The excitation functions of the Au(,n)Au and Au(,3n)Au reactions were determined in the incident photon energy range of 13-31 MeV using quasi-monoenergetic photon beams provided at HIS. The cross sections of the combined ground state (2) and short-lived first isomeric state (m1, 5), and of the second isomeric state (m2, 12) in the Au production are obtained separately by subtracting the rays from the internal conversion of the second isomeric state. The excitation function of the second isomeric state via the photon-induced reaction Au(,n)Au was measured for the first time. By using the activation method rather than direct neutron counting, the exclusive cross sections for the (,n) and (,3n) reactions were determined. Comparing the yields from the front and back gold targets validates our ability to simulate the effect of photon scattering in the target stack and provides a method for assessing the systematic uncertainty of our technique.
I Introduction
Radioisotopes play important roles in numerous fields ranging from medical treatments to national security and basic research. Some examples include investigations of structures and reactions involving atomic nuclei, Mossbauer spectroscopy, radio-thermoelectric generation and other nuclear batteries, nuclear device detection, and the mitigation of nuclear proliferation [1]. The reports from “Workshop on the Nation’s Needs for Isotopes: Present and Future”, “Isotopes for the Nation’s Future, A long range plan”, and “Meeting Isotope Needs and Capturing Opportunities for the Future” identify multiple isotopes in short supply [1, 2, 3]. Among these are high specific activity beta emitters such as Sc, Cu, As, and Re. These radioisotopes are of interest to the community as they have ideal nuclear properties for medical applications. Their availability is limited as they are difficult to produce in high specific activity using common production methods.
Most radioisotopes are produced using either nuclear reactors or light-ion accelerators. High specific activity radioisotopes are very important in nearly all aspects of radiochemical work. However, most radionuclides produced in nuclear reactors have low specific activity due to the inherently limited production pathways, e.g., dominated by (n,) reactions. Specific activity refers to the radioactivity of the produced radionuclide relative to the total mass of the element of interest present in the sample. In contrast, accelerator-produced radionuclides are generally characterized by high specific activity, e.g., (p,xn) reactions are typically used because the produced isotope is a different element from the target material. Electron linear accelerators (LINACs) are unique sources of radioisotopes. Even though the basic technology has been around for decades, only recently have electron LINACs capable of producing photons with sufficient energy and flux for radioisotope production become available. The Nuclear Science Advisory Committee (NSAC) recently named this approach as one of the most compelling with the largest-impact opportunities for the production of high specific activity radioisotopes [1]. Electron LINACs activate materials via photonuclear reactions. The yields of these reactions depend on the production of high-energy photons generated by interactions of high-energy electrons with a high-Z material (i.e., a converter) to produce Bremsstrahlung radiation. Production yields are controlled by the electron beam energy, beam flux, target size, length of irradiation, and reaction cross section. A great deal of work has been done to create experimental cross-section databases for (,n) reactions (leading towards low specific activity radioisotopes) [4, 5, 6]. However, these data are incomplete. For example, the cross-section data for (,p) reactions (which lead to high specific activity radioisotopes) still require considerable work. There are many theoretical databases available which provide guidance on photo-production of radionuclides. Experimental data are needed to assess the accuracy of predictions of the leading reaction models used in nuclear-data libraries (TENDL, JENDL, ENDF, JEFF, CENDL, BROND, etc.) [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Examples of theoretically predicted cross sections that incorrectly estimate the production rate of radioisotopes are available in the literature [17, 18, 19, 20]. One specific example is the photonuclear production of Cu via the reaction Zn(,p)Cu where the maximum theoretical cross section in the TENDL database is 3.7 mb at 22 MeV, but was experimentally determined to be 11.76 mb as reported by the Experimental Nuclear Reaction Data (EXFOR) database. This result effectively implies that Cu can be produced more economically via the (,p) reaction than previously thought. In order to better assess the feasibility of producing Cu and other radioisotopes with photonuclear reactions, (,n) or (,p), excitation functions must be experimentally evaluated. Direct measurement of photonuclear cross sections using electron beams is not trivial as converter-produced Bremsstrahlung photons deliver a beam with a broad energy spectrum to the targets [21, 22]. Compton back-scattering is an elegant solution to this problem. Compared with a Bremsstrahlung beam, with a broad-band energy continuum, a Compton back-scattered -ray beam is narrowly peaked around a specific energy. Very precise measurement of cross sections can be obtained using activation techniques with target irradiation by these quasi-monochromatic gamma beams. Such quasi-monoenergetic photon beams are available at the Triangle Universities Nuclear Laboratory’s (TUNL) High-Intensity Gamma-Ray Source (HIS) at the Duke Free-Electron-Laser Laboratory (DFELL). Photons with selectable energy are produced using intracavity backscattering of free-electron laser (FEL) photons from electrons in the DFELL storage ring. Gamma-ray beams with photon energies of up to 100 MeV and total intensities of 10 - 10 /s can be produced [23]. High-energy photons are produced by colliding electron bunches that are synchronized so that laser photons from one bunch are reflected from a downstream mirror and then collide head-on with the second electron bunch in a magnetic-field-free scattering region. Photon beam energies and fluxes are calibrated with in-line detectors and activation of well-defined monitor foils, respectively. Additionally, the back-scattered gamma beam has a well-defined radial energy dependence. The central portion of the beam has the highest energy. The beam energy and flux density decrease radially outward from the center in a predictable manner. The beam flux also surrenders to this phenomenon [24]. Activation of concentric ring targets by an uncollimated beam will provide experimental activation data that may be used to calculate cross sections at the energy of the beam incident on each ring. This method was developed and used for the first time at HIS in the present energy-dependent cross-section measurements which provide activation data at 3 energies for each target irradiation.
In this paper, we describe our newly developed experimental techniques for measuring photonuclear cross sections using the laser-Compton scattered beam at HIS, and report cross sections for photon-induced reactions on Au measured using these new techniques. The cross section of Au(,xn) reactions are important for beam flux measurements in cross section experiments. For example, gold monitor foils are routinely used to normalize Bremsstrahlung photon beams in cross-section measurements [25, 26]. The Au(,n)Au cross section is very well characterized in the energy range of 8-20 MeV, which is below the threshold for multiple neutron emission (see Table 1). However, there are discrepancies between existing experimental data [27, 28], and above 20 MeV gamma energy, there are very few experimental data and theoretical predictions by reaction models differ substantially. These photon energies are above the threshold for reactions leading to high specific activity radionuclides via the (,p) reaction, such as Ti(,p)Sc and Zn(,p)Cu at around 20 MeV and higher. A goal of the present work is to measure the exclusive cross-section excitation functions of the Au(,n) and (,3n) reactions with improved accuracy over existing data for use as a standard in determining -ray beam fluxes. Table 1 provides a summary of the world data for the Au(,xn) reactions by beam type. These data are compared to the results of our new measurements in Section V. The results of this work for the Ti, Zn, and Os targets measured concurrently with Au will be presented in a subsequent paper. In this paper, our multiple target multi-energy technique is described in section II, the HIS Laser Compton Scattering (LCS) beam is described in section III, the experiment is discussed in section IV, section V describes the cross-section analysis, and the results are presented and discussed in section VI.
| Reaction | Reference | Year | range (MeV) | Points | Lab | Beam type | Technique |
| (,n) | C. Plaisir et al. [27] | 2012 | 10-20 | 101 | CEA/DAM ELSA, France | Bremsstrahlung | activation counting |
| (,n) | K. Vogt et al. [29] | 2002 | 8.08-9.90 | 9 | TU-Darmstadt, Germany | Bremsstrahlung | activation counting |
| (,n) | B.L. Berman et al. [30] | 1987 | 12.1-16.9 | 11 | LLNL, USA | annihilation | neutron counting |
| (,n) | A. Veyssiere et al. [31] | 1970 | 8.08-19.8 | 43 | CEA Paris-Saclay, France | annihilation | neutron counting |
| (,n) | S.C. Fultz et al. [28] | 1962 | 8.7-17.9 | 30 | AIST, Tsukuba, Japan | annihilation | neutron counting |
| (,n) | O. Itoh et al. [32] | 2011 | 8.08-13.14 | 30 | AIST, Tsukuba, Japan | LCS source | neutron counting |
| (,n) | F. Kitatani et al. [33] | 2011 | 9.3-12.4 | 12 | AIST, Tsukuba, Japan | LCS source | neutron counting |
| (,n) | F. Kitatani et al. [34] | 2010 | 10.0-14.6 | 8 | AIST, Tsukuba, Japan | LCS source | neutron counting |
| (,n) | K.Y. Hara et al. [35] | 2007 | 8.23-12.3 | 24 | AIST, Tsukuba, Japan | LCS source | neutron counting |
| (,2n) | B.L. Berman et al. [30] | 1987 | 14.0-16.9 | 7 | LLNL, USA | annihilation | neutron counting |
| (,2n) | A. Veyssiere et al. [31] | 1970 | 13.5-27.1 | 40 | CEA Paris-Saclay, France | annihilation | neutron counting |
| (,2n) | S.C. Fultz et al. [28] | 1962 | 14.3-24.7 | 29 | LLNL, USA | annihilation | neutron counting |
| (,3n) | A. Veyssiere et al. [31] | 1970 | 22.8-27.1 | 8 | CEA Paris-Saclay, France | annihilation | neutron counting |
II Method
This section details the method applied in this work to determine the exclusive (,n) and (,3n) cross sections via the activation method. The excitation functions reported here cover the gamma-energy range from 13 to 31 MeV. The photonuclear cross sections over the range of the Giant Dipole Resonance (GDR) vary by orders of magnitude. The goal of this work is to demonstrate our activation techniques by applying them in measuring the excitation functions of Au(,xn) reactions with a resolution of 3-10 at each energy. At HIS, the beam energy spread depends on the beam collimation. In a typical setup with an energy spread (FWHM) of 5, the flux transmitted through the collimator is about 7.5 of the total flux of the Compton-scattered laser beam. Our method of using 3 radial target segments enables 3 gamma energies of the excitation function to be measured concurrently. The radial energy and flux distribution of the HIS beam was verified experimentally by blocking portions of the beam with tungsten plugs to mimic the concentric-rings target design. Simulations were performed and compared to the experimental results of these calibrations to determine the accuracy of the beam simulations.
II.1 Target stack
In the activation method, the photon beam induces radioactivity via photon-induced reactions, such as (,n) and (,p), in targets. Reaction yields are determined by gamma counting the targets after activation. The target foils used in these experiments are made of separable concentric rings that exploit the energy spread of the HIS beam and enable the activation of multiple targets across multiple energies in a single beam exposure. A concept of a target stack is illustrated in Figure 1. The relatively low and well characterized attenuation of high-energy photons by the target materials provides the ability to irradiate multiple targets in a single irradiation. The total thickness of the stack of targets used in these experiments was 2.8-3.3 g/cm, which resulted in a 13 total attenuation of the photon beam at 14 MeV. The entire target may then be exposed to the beam and each individual concentric ring will interact with a narrow range of beam energies and a calculable fraction of the overall photon flux. The individual rings are analyzed with high-purity germanium detectors (HPGe) to determine activation yields. Thus, using this technique, the cross sections at 3 energies can be measured in a single irradiation. With the stack of 5 targets and 3 segments per target, yields for 15 target segments are obtained for each 12-hour run.
II.2 Activation
The reaction rate (R) is given by,
where is the cross section, is the number of atoms in the target, and I is the beam flux of incident particles.
An example of a radioisotope production curve is shown in Figure 2. The isotope is being produced faster than it can decay during the time from 0 to when the target is being irradiated.
II.2.1 Au(,n)Au
The Au nucleus has three different long-lived states; e.g. the ground, and isomeric states denoted by g (6.17 d), (8.1 s) and (9.6 h). The two isomeric states decay 100 to the ground state. The number of nuclei produced during the irradiation time can be calculated using the formulae below,
where the R’s and ’s are the reaction rates and decay constants, respectively. is the number of nuclei in the th state, here numbers 0,1 and 2 represent ground, isomeric and isomeric state, respectively. The solutions for the number of nuclei in the and isomeric states are obtained from the following equations.
The number can be approximately constant, since due to the extremely short half-life compared to that of ground and isomeric states. This means . In other words, we can determine the cross sections for the production of and , separately. Using the above solutions and equation 2, we find the solution for ,
where . When the irradiation is finished, the nuclei will decay with no production term. The number of nuclei decreases with the respective half-lives,
The number of nuclei can be calculated from integrating equation 3,
with,
The total number of events with measurement time() will be,
Using these results and the relation between the number of events and the cross section in equation 1, the cross section for producing Au in the 2 isomer state is,
Similarly, the combined cross section for ground and isomeric state can be obtained,
with,
II.2.2
The cross section for producing is determined in the same way as Equation 4,
III quasi-monoenergetic Photon beam
The photon beams are produced by Compton scattering of FEL photons from relativistic electrons circulating in a storage ring at HIS. In the case of relativistic electrons ( 1) scattering at a small angle (1), the scattered photon energy after collision is given by,
where E is the scattered photon energy, and E and E are the energies of the incident photon and electron. The Lorentz factor of the electron is denoted as . The angle is expressed in terms of the radius of the target (r) and the distance from the collision point to the target position, r/L. The spatial distribution of the scattered photons can be obtained theoretically [24]. The theoretical predictions are verified using a computer code [24] developed by HIS’s accelerator group, which was benchmarked with other codes such as [36] and using experimental results at HIS [24]. As an example of Monte-Carlo simulation results based on the theoretical prediction, if we consider the -ray beam at HIS with E 660 MeV and 450 nm, the scattered -ray beam energies at the target location (L 53 m) at each of the three segments will be 18, 17 and 16 MeV, respectively. Typical percentile energy spreads (1) of the photon beam at the 1, 2 and 3 segments are 1.3-2.7, 2.2-4.3 and 3.2-5.7 of the central beam energy from 14 to 31 MeV, respectively. The spatial distributions and photon energies in each target segment from simulations are provided in Figure 3, respectively. The accuracy of the simulations of the beam energy profile in different rings concentric with the beam axis have been validated by measurements. These measurements were performed using rays produced under conditions identical to those used in the simulations, E 725 MeV and 350 nm. The energy spectrum of the beam was measured using a large cylindrical NaI detector (10” dia. 14” thick) positioned with its center on the -ray beam axis. This detector is referred to as “Molly”. Data were taken for four collimator arrangements: N1 = 0.0 to 6.35 mm; N2 = 0.0 to 15.9 mm; N3 = 6.35 to 15.9 mm; and N4 = 11.1 to 15.9 mm. The efficiency of the large NaI detector (Molly) is greater than 97 at the energies of these measurements. The N1 and N2 measurements were performed using standard lead collimators with openings of 0.5-inch and 1.25-inch diameter. The N3 and N4 measurements were carried out by inserting a 0.5-inch and a 0.875-inch diameter tungsten rod concentric with the 1.25-in diameter collimator, respectively. Both tungsten rods were 6 inches long.
The beam flux at each segment of each target is essential in determining the corresponding reaction cross sections. The fractional photon beam fluxes at each target segment were obtained from simulations. With these incident beam fluxes on the segments of the front target, the beam intensity at each target was calculated using the beam transport simulation performed by the PHITS Monte-Carlo code [37]. The geometry illustrated in Figure 4 was used for the simulations. The targets in the stack were Au, TiO, Zn, Os and Au. Successive targets were separated by 0.45 cm. The detailed information about the segments of the actual target stack can be found in Section IV. A total of 15 segments in the target stack (3 segments per target) were exposed to the photon beams at once. The beam flux was simulated for each irradiation (See Table 2). In this energy range, Compton scattering and pair production dominate the beam attenuation and scattering as it passes through the stack. The attenuation and scattering of the gamma beam is simulated in detail using the PHITS code with the overall attenuation predicted to be 10. The central beam energy, one-sigma radial energy spread, and beam flux on each segment of the stack are summarized in Table 2. The overall simulation is validated using the data from the gold target at the back of the stack when compared with the data from the gold target in the front.
| E | 14 MeV | 16 MeV | 18 MeV | 22 MeV | |||||||||
| seg | tgt. | flux () | flux () | flux () | flux () | ||||||||
| 1 | Au | 13.956 | 0.181 | 23.960 | 15.944 | 0.227 | 24.400 | 17.907 | 0.243 | 24.920 | 21.828 | 0.483 | 24.680 |
| TiO | 13.956 | 0.182 | 23.080 | 15.944 | 0.226 | 23.430 | 17.906 | 0.243 | 23.840 | 21.825 | 0.497 | 23.536 | |
| Zn | 13.956 | 0.181 | 22.960 | 15.944 | 0.226 | 23.300 | 17.906 | 0.243 | 23.710 | 21.824 | 0.500 | 23.404 | |
| Os | 13.956 | 0.182 | 22.670 | 15.944 | 0.227 | 23.090 | 17.906 | 0.243 | 23.520 | 21.823 | 0.504 | 23.204 | |
| Au | 13.956 | 0.182 | 22.300 | 15.944 | 0.227 | 22.720 | 17.906 | 0.243 | 23.050 | 21.822 | 0.510 | 22.709 | |
| 2 | Au | 13.477 | 0.303 | 23.320 | 15.303 | 0.430 | 23.180 | 17.100 | 0.486 | 23.160 | 20.903 | 0.692 | 23.285 |
| TiO | 13.477 | 0.303 | 22.280 | 15.303 | 0.430 | 21.990 | 17.100 | 0.486 | 22.300 | 20.901 | 0.697 | 22.363 | |
| Zn | 13.477 | 0.303 | 22.180 | 15.302 | 0.430 | 21.890 | 17.100 | 0.486 | 22.220 | 20.901 | 0.697 | 22.276 | |
| Os | 13.477 | 0.303 | 21.960 | 15.302 | 0.430 | 21.730 | 17.100 | 0.486 | 22.000 | 20.900 | 0.699 | 22.053 | |
| Au | 13.477 | 0.303 | 21.650 | 15.302 | 0.430 | 21.440 | 17.100 | 0.486 | 21.730 | 20.900 | 0.700 | 21.763 | |
| 3 | Au | 12.915 | 0.414 | 19.270 | 14.581 | 0.468 | 18.830 | 16.184 | 0.637 | 18.610 | 19.854 | 0.824 | 18.871 |
| TiO | 12.914 | 0.414 | 18.230 | 14.581 | 0.468 | 17.880 | 16.184 | 0.637 | 17.720 | 19.852 | 0.827 | 17.903 | |
| Zn | 12.914 | 0.414 | 18.140 | 14.581 | 0.468 | 17.780 | 16.184 | 0.637 | 17.620 | 19.852 | 0.828 | 17.807 | |
| Os | 12.914 | 0.414 | 17.980 | 14.581 | 0.468 | 17.630 | 16.183 | 0.637 | 17.430 | 19.851 | 0.828 | 17.609 | |
| Au | 12.914 | 0.415 | 17.410 | 14.580 | 0.468 | 17.100 | 16.183 | 0.638 | 17.090 | 19.851 | 0.829 | 17.235 | |
| E | 24 MeV | 26 MeV | 31 MeV | ||||||||||
| seg | tgt. | flux () | flux () | flux () | |||||||||
| 1 | Au | 24.115 | 0.573 | 24.619 | 26.232 | 0.655 | 25.179 | 31.053 | 0.843 | 26.180 | |||
| TiO | 24.111 | 0.596 | 23.525 | 26.227 | 0.688 | 24.089 | 31.040 | 0.933 | 24.850 | ||||
| Zn | 24.110 | 0.602 | 23.389 | 26.225 | 0.697 | 23.958 | 31.037 | 0.948 | 24.710 | ||||
| Os | 24.109 | 0.609 | 23.159 | 26.222 | 0.714 | 23.620 | 31.034 | 0.968 | 24.491 | ||||
| Au | 24.107 | 0.619 | 22.741 | 26.219 | 0.732 | 23.168 | 31.007 | 1.125 | 23.068 | ||||
| 2 | Au | 22.987 | 0.830 | 23.308 | 24.912 | 0.948 | 23.239 | 29.245 | 1.264 | 23.080 | |||
| TiO | 22.982 | 0.844 | 21.974 | 24.906 | 0.969 | 22.019 | 29.237 | 1.295 | 22.083 | ||||
| Zn | 22.982 | 0.845 | 21.870 | 24.906 | 0.970 | 21.909 | 29.237 | 1.297 | 21.994 | ||||
| Os | 22.981 | 0.846 | 21.691 | 24.905 | 0.973 | 21.665 | 29.234 | 1.305 | 21.763 | ||||
| Au | 22.980 | 0.847 | 21.369 | 24.904 | 0.977 | 21.316 | 29.220 | 1.362 | 20.687 | ||||
| 3 | Au | 21.718 | 0.986 | 18.757 | 23.431 | 1.131 | 18.404 | 27.239 | 1.509 | 17.814 | |||
| TiO | 21.714 | 0.993 | 17.694 | 23.425 | 1.144 | 17.238 | 27.230 | 1.532 | 16.818 | ||||
| Zn | 21.714 | 0.993 | 17.599 | 23.425 | 1.144 | 17.150 | 27.229 | 1.533 | 16.726 | ||||
| Os | 21.714 | 0.994 | 17.433 | 23.424 | 1.145 | 16.980 | 27.228 | 1.535 | 16.530 | ||||
| Au | 21.712 | 0.997 | 16.857 | 23.421 | 1.151 | 16.347 | 27.217 | 1.563 | 15.582 |
IV Experiment
Targets were activated at seven different central photon beam energies (see Table 3). The gamma energy at the outermost ring typically was close to the central energy of the next lower energy irradiation, thus providing an internal consistency check for these measurements. Using targets that of three concentric segments, the cross-section measurements for the reactions Au(,n)Au and Au(,3n)Au were determined for 21 different energies. The target stack was irradiated for 8-12 hours by a photon beam with fluxes that ranged between 10-10 /s, limited by a copper pre-collimator. The boundaries of the photon beam of 30-40 mm in diameter were shaped by the copper apertures inside the FEL aperture systems and by the copper pre-collimator. Irradiation times, beam intensities and uncertainties are summarized in Table 3.
| Irradiation time | Intensity | Unc. | |
| MeV | hr | /s | |
| 14 | 11.49 | 8.01 | 11 |
| 16 | 10.72 | 8.80 | 11 |
| 18 | 8.17 | 9.27 | 11 |
| 22 | 10.73 | 2.48 | 11 |
| 24 | 11.20 | 2.70 | 11 |
| 27 | 11.73 | 2.51 | 11 |
| 31 | 9.12 | 2.06 | 11 |
IV.1 Target stack
The target stack was composed of layers of Au, TiO, Zn, Os and Au as shown in Figure 5 and listed in Table 4. The target segments were mounted in 3D-printed acrylic with cavities made for receiving powdered target material. The cavities were filled with a known mass of the target material and then sealed by adding and curing 3D printing gel on top of the target material. The effects of photon attenuation and scattering by the low-acrylic target mounts were negligible. The thickness of the targets varied across the numerous targets made for this work (see Table 4). The front and back Au targets were arranged in this manner in anticipation of using these to normalize the beam flux and calibrate the beam energy throughout the stack. Placed front to back in the stack, relative to the incident beam, and between the front and back Au targets, were TiO, Zn and Os ones. The beam flux and scattered gamma contribution to the cross section were calculated. There were 3 sets of targets to enable reuse after sufficient decay of the activations in subsequent runs separated by months.
| 14, 27 | 16, 24 | 18, 22, 31 | |||
| thick. | |||||
| tgt. | seg. | (g/cm) | (g/cm) | (g/cm) | |
| 1 | Au | 1 | 0.681 | 0.713 | 0.763 |
| 2 | 0.931 | 0.933 | 0.655 | ||
| 3 | 1.033 | 0.935 | 0.869 | ||
| 2 | TiO | 1 | 0.228 | 0.241 | 0.235 |
| 2 | 0.205 | 0.194 | 0.163 | ||
| 3 | 0.205 | 0.220 | 0.221 | ||
| 3 | Zn | 1 | 0.501 | 0.350 | 0.308 |
| 2 | 0.392 | 0.292 | 0.363 | ||
| 3 | 0.351 | 0.339 | 0.406 | ||
| 4 | Os | 1 | 0.304 | 0.293 | 0.358 |
| 2 | 0.257 | 0.242 | 0.220 | ||
| 3 | 0.614 | 0.553 | 0.362 | ||
| 5 | Au | 1 | 1.209 | 1.299 | 1.115 |
| 2 | 0.834 | 1.195 | 1.049 | ||
| 3 | 1.304 | 1.395 | 1.471 |
IV.2 Photon beam flux
A significant portion of the overall systematic uncertainty in the cross sections measured in this work is due to the errors in measuring the absolute -ray beam flux. A schematic diagram of the system used to measure this flux during target irradiation is given in Figure 6. The beam flux was measured using three detectors during target irradiation: 1) a shielded plastic scintillator paddle that detects rays and recoil electrons from Compton scattering of the beam passing through the mirror of the optical cavity (labeled MP in Figure 6), 2) a thin plastic scintillator centered in the beam path just after the collimator (labeled SP in Figure 6), and 3) a silicon detector that measures charged particles emitted from photon-induced reactions in a thin CD foil inside an evacuated scattering chamber (labeled CP in Figure 6). The MP detector monitors a portion of the full -ray beam intensity passing through the cavity mirror. The detection efficiency of MP for measuring the beam passing through the collimator is determined using Molly (see Figure 6). This detector can be positioned in the beam path, and a system of high-precision copper absorbers can be inserted into the beam to attenuate the incident flux on Molly to keep the data acquisition system deadtime lower than 40 and the pulse pileup rate of the detector signal below 10. The efficiencies of the other flux monitors, SP and CP, are calibrated relative to the MP using the full beam intensity. The efficiency of MP is determined using the following relationships,
where . The symbol is the photon attenuation coefficient (cm/g) averaged over the beam intensity profile, is the density of the copper attenuators (8.96 g/cm), and x is the length in cm of the copper attenuator inserted in the beam path to reduce the flux incident on Molly. The value of was calculated using the photon attenuation coefficients as a function of energy, (E), obtained from NIST [38, 39], and the beam intensity profile simulated as described in [40]. An example of and , the unattenuated and attenuated beam intensity profiles, used to calculate at the peak beam energy, = 18.0 MeV, is plotted in Figure 7. The value is the -ray energy, E, where the distribution has a maximum value. The distribution is normalized such that the integration over the distribution is equal to 1.0. Because the uncertainty in has a significant impact on the error of the computed -ray transmission through the copper absorbers, the calculated value was benchmarked at several beam energies with in-situ measurements. The average variance between the calculated and measured values was 1.0. This variance was assigned as an estimate of the uncertainty in the calculated values. In turn, the uncertainty in was used to compute the average error in the calculated intensity. The main systematic uncertainty in the measurement of the -ray beam flux arises from the error in determining the effective x-ray attenuation coefficient, , used to compute the -ray transmission through the copper absorbers. The systematic uncertainty in the -ray beam flux determination is 11.0.
IV.3 -ray spectra
After activation, -ray spectra were measured using 12 High Purity Germanium (HPGe) detectors, each with a co-axial cylindrical crystal positioned with its front face 5 cm from the activated target. The measurements started approximately 30 minutes after the end of bombardment and lasted long enough to obtain sufficient counts for the second isomeric state of Au with the half-life of 9.6 hours. Each HPGe detector was calibrated with a mixed calibration source, and the detector efficiencies for the gold isotopes were determined by the given formula [42].
where x and y are ln and ln/1000), respectively. E is the -ray energy from selected nuclei. The absolute efficiency at 333 keV for the Au transition and the corresponding uncertainty is 1-3 and 2, depending on the HPGe being considered. During the measurements, the dead time was typically 1-2. The spectra at the photon beam energies of 18 and 27 MeV are presented in Figure 8. The emitted -ray peaks of 147.81, 188.27 keV (second isomeric state of Au), 333.03, 355.73 keV (ground state of Au) and 293.55, 328.47 keV (Au) are visible in Figure 8.
V Cross-section analysis
The photonuclear cross sections were determined using equations 4 and 5. In addition, the detector efficiency, detector dead time, branching ratio, beam attenuation and self-absorption effects are taken into account. With these corrections, the cross sections can be expressed as,
where and are the cross sections for the ground + 1 isomeric state and 2 isomeric state, respectively. stands for the detected counts, and are the beam flux and the detector efficiency, respectively. The total correction factor above is expressed as = . The , , , and symbols are corrections for the detector efficiency, detector dead time during the measurements, -ray branching ratio, -ray beam attenuation and self-absorption by passing through the targets, respectively. Correction of the photon beam flux due to attenuation, and of the emitted -ray yield due to self-absorption are taken into account by,
where and x are attenuation coefficients for the beam and emitted -ray and the total length of the target, respectively. The statistical error was calculated by RADWARE software [42] and three systematic errors for the -ray branching ratio, detector efficiency and beam flux were determined. Therefore, the total experimental error can be obtained by,
where , , and are uncertainties from the number of nuclei measured by counts, photon beam flux, HPGe detector efficiency and the branching ratio of the selected nuclei, respectively.
After activation by the photon beam, the segments in the target stacks were separated and counted individually with HPGe detectors. The perceptible Au radioisotopes produced via Au(,xn) were Au and Au. The Au nucleus was observed to produce two isomeric states (m= 5, m= 12) and the ground state (2). The simplified decay schemes for the Au and Au nuclei are illustrated in Figure 9. The half-life of the , and ground states of Au are 8.1 s, 9.6 h and 6.17 d, respectively. The m2 and ground states are measurable and distinguishable by counting the rays from the gold targets within a few hours from the end of each irradiation period. The m isomeric state, with its short half-life of only 8.1 s, has its yield included in the decay of the ground state. The 147.81 and 188.24 keV -ray lines shown in Figure 9 and the peaks in Figure 8 are used to determine the yield of the m isomer, and the 333.03, 328.47 keV -ray transitions following emission from ground state of Au are used to calculate the cross section for the sum of the ground state and m1 isomer, respectively. The ground state (1) of Au has a 38.02 h half-life and decays by emission to Pt. The 293.55 and 328.47 keV -ray peaks were used to determine the cross section of the Au(,3n)Au reaction. The decay information for the Au and Au nuclei, along with uncertainties, is listed in Table 5.
| nucl. | half-life | peak (keV) | branch ratio () |
| Au | 6.1669 (6) d | 333.03 | 22.9 (9) |
| 355.73 | 87 (3) | ||
| Au | 9.6 (1) h | 147.81 | 43.5 (15) |
| 188.27 | 30.0 (15) | ||
| Au | 38.02 (10) h | 293.549 | 10.9 (3) |
| 328.470 | 62.8 (16) |
V.1 Au(,n)Au
The cross sections of the combined ground and first isomeric states and second isomeric state for the photon-induced reaction on Au are obtained with the emitted rays of 333.03, 355.73 keV and 147.81, 188.27 keV, respectively. The excitation functions for the Au(,n)Au and Au(,n)Au reactions are presented in Figure 10 together with the previous experimental data [27, 28, 35, 34, 32, 33, 31, 30, 29] and theoretical calculations from TENDL 2021 and JENDL 4.0. Here, the cross sections in Figure 10 were determined by means of weighted mean and error. The cross sections measured in this work with statistical and total uncertainties are given in Table 6.
| nucl. | Au | Au | |||||||||||||||
| peak | 333.03 | keV | 355.73 | keV | 147.81 | keV | 188.27 | keV | |||||||||
| stat. | tot. | stat. | tot. | stat. | tot. | stat. | tot. | ||||||||||
| MeV | MeV | mb | mb | mb | mb | mb | mb | ||||||||||
| 12.95 | 0.41 | 518.3 | 0.2 | 11.9 | 512.3 | 0.1 | 11.7 | 515.0 | 11.4 | 0.16 | 63.6 | 64.7 | 0.16 | 50.4 | 51.9 | 0.16 | 41.2 |
| 13.48 | 0.30 | 545.3 | 0.2 | 11.9 | 541.7 | 0.1 | 11.7 | 543.3 | 11.4 | 0.24 | 2.3 | 11.9 | 0.25 | 38.4 | 40.3 | 0.24 | 11.9 |
| 13.96 | 0.18 | 533.0 | 0.2 | 11.9 | 530.8 | 0.1 | 11.7 | 531.8 | 11.4 | 0.28 | 16.9 | 20.6 | 0.27 | 9.2 | 15.3 | 0.27 | 14.3 |
| 14.58 | 0.47 | 500.9 | 0.2 | 11.9 | 496.4 | 0.1 | 11.7 | 498.4 | 11.4 | 0.37 | 5.3 | 12.8 | 0.35 | 9.9 | 15.8 | 0.36 | 12.4 |
| 15.30 | 0.43 | 370.5 | 0.2 | 11.9 | 367.1 | 0.1 | 11.7 | 368.6 | 11.4 | 0.41 | 8.3 | 14.3 | 0.44 | 5.8 | 13.5 | 0.42 | 12.5 |
| 15.94 | 0.23 | 241.5 | 0.3 | 11.9 | 240.4 | 0.2 | 11.7 | 240.9 | 11.4 | 0.39 | 6.1 | 13.2 | 0.42 | 6.4 | 13.8 | 0.40 | 12.3 |
| 16.18 | 0.64 | 193.7 | 0.4 | 11.9 | 192.6 | 0.2 | 11.7 | 193.1 | 11.4 | 0.30 | 7.7 | 14.0 | 0.33 | 6.6 | 13.9 | 0.31 | 12.6 |
| 17.10 | 0.49 | 93.1 | 0.5 | 11.9 | 92.2 | 0.3 | 11.7 | 92.6 | 11.4 | 0.20 | 5.8 | 13.1 | 0.25 | 6.4 | 13.8 | 0.22 | 12.2 |
| 17.91 | 0.24 | 54.7 | 0.6 | 11.9 | 55.3 | 0.3 | 11.7 | 55.0 | 11.4 | 0.20 | 6.1 | 13.2 | 0.19 | 5.7 | 13.5 | 0.20 | 12.2 |
| 19.85 | 0.82 | 20.3 | 2.0 | 12.0 | 21.1 | 1.1 | 11.8 | 20.7 | 11.4 | 0.12 | 14.0 | 18.3 | 0.13 | 14.1 | 18.7 | 0.12 | 15.2 |
| 20.90 | 0.69 | 15.3 | 4.1 | 12.5 | 14.7 | 0.7 | 11.7 | 14.9 | 11.5 | 0.12 | 9.3 | 14.9 | 0.11 | 17.4 | 21.3 | 0.12 | 14.1 |
| 21.72 | 0.99 | 12.7 | 2.7 | 12.4 | 12.2 | 1.6 | 11.8 | 12.4 | 11.5 | 0.10 | 11.0 | 13.6 | 0.09 | 13.7 | 16.7 | 0.09 | 12.9 |
| 21.83 | 0.48 | 13.6 | 3.5 | 12.2 | 12.6 | 1.2 | 11.8 | 13.0 | 11.5 | 0.10 | 7.0 | 16.1 | 0.10 | 11.3 | 18.4 | 0.10 | 14.3 |
| 22.99 | 0.83 | 11.2 | 2.7 | 12.2 | 10.4 | 4.5 | 11.9 | 10.7 | 11.5 | 0.09 | 16.0 | 19.9 | 0.09 | 7.1 | 14.1 | 0.09 | 13.5 |
| 23.43 | 1.13 | 9.3 | 7.2 | 12.6 | 9.3 | 2.8 | 11.8 | 9.3 | 11.5 | 0.09 | 12.4 | 44.8 | 0.08 | 4.5 | 26.6 | 0.08 | 23.8 |
| 24.12 | 0.57 | 8.3 | 4.2 | 13.9 | 8.5 | 1.4 | 12.0 | 8.5 | 11.8 | 0.09 | 43.2 | 17.0 | 0.08 | 23.6 | 13.0 | 0.08 | 12.6 |
| 24.91 | 0.95 | 8.0 | 3.3 | 12.3 | 8.2 | 1.4 | 11.8 | 8.1 | 11.5 | 0.09 | 43.1 | 44.7 | 0.08 | 15.2 | 19.5 | 0.08 | 18.7 |
| 26.23 | 0.66 | 5.7 | 9.0 | 14.9 | 6.2 | 7.4 | 13.9 | 6.0 | 12.8 | 0.08 | 32.6 | 34.6 | 0.07 | 16.1 | 20.2 | 0.07 | 18.6 |
| 27.24 | 1.51 | 5.2 | 2.6 | 12.1 | 4.3 | 0.5 | 11.7 | 4.6 | 11.4 | 0.08 | 19.8 | 23.0 | 0.06 | 29.7 | 32.1 | 0.07 | 20.1 |
| 29.25 | 1.26 | 3.8 | 3.5 | 12.4 | 3.9 | 1.7 | 11.8 | 3.9 | 11.5 | 0.07 | 41.5 | 43.1 | 0.07 | 20.1 | 23.6 | 0.07 | 21.6 |
| 31.05 | 0.84 | 3.2 | 4.4 | 12.6 | 3.0 | 1.7 | 11.8 | 3.1 | 11.6 | 0.05 | 18.3 | 21.7 | 0.05 | 17.3 | 21.2 | 0.05 | 17.0 |
V.2 Au(,3n)Au
The cross sections for the Au(,3n)Au reaction are obtained from the emitted rays of 293.55 and 328.47 keV. The weighted mean cross sections are presented in Figure 11, together with the previous experimental data [31] and theoretical calculations from TENDL 2021 and JENDL 4.0. The cross sections measured in this work with statistical and total uncertainties are summarized in Table 7.
| nucl. | Au | ||||||||
| peak | 293.55 | keV | 328.46 | keV | |||||
| stat. | tot. | stat. | tot. | ||||||
| MeV | MeV | mb | mb | mb | |||||
| 22.99 | 0.83 | - | - | - | 0.12 | 5.31 | 13.6 | 0.12 | 13.6 |
| 23.43 | 1.13 | 0.86 | 6.3 | 12.9 | 0.90 | 2.13 | 11.5 | 0.89 | 11.4 |
| 24.12 | 0.57 | 1.83 | 7.6 | 11.6 | 2.00 | 1.22 | 11.3 | 1.93 | 11.2 |
| 24.91 | 0.95 | 7.63 | 2.2 | 11.5 | 8.02 | 0.92 | 11.3 | 7.86 | 11.2 |
| 26.23 | 0.66 | 16.18 | 0.9 | 11.3 | 17.02 | 1.91 | 11.4 | 16.51 | 11.2 |
| 27.24 | 1.51 | 16.24 | 1.7 | 11.4 | 15.78 | 0.44 | 11.3 | 15.96 | 11.2 |
| 29.25 | 1.26 | 21.25 | 1.2 | 11.3 | 24.54 | 0.39 | 11.3 | 22.86 | 11.1 |
| 31.05 | 0.84 | 20.41 | 1.3 | 11.3 | 21.70 | 0.41 | 11.3 | 21.10 | 11.2 |
VI Results and Discussion
This work demonstrates the successful development of a novel and highly efficient method for concurrently measuring photonuclear reaction cross sections at multiple discrete photon energies for various materials using a single irradiation. As a proof-of-concept, the excitation function of the Au(,n)Au and Au(,3n)Au reactions were measured with this novel method utilizing target stacks consisting of concentric ring segments that were bombarded with an uncollimated beam with well-defined radial energy and flux distributions. The measurements were performed at 21 discrete beam energies over the incident photon beam energy range of 13-31 MeV (see Table 6). For the first time, the cross sections of the combined ground and first isomeric states at the beam energy above 20 MeV and the second isomeric state for the Au(,n)Au reactions were determined via the activation method which yielded exclusive cross sections which are unaffected by multiple neutron emission. The method developed here based on radial target segmentation and target stacking makes efficient use of beam time by concurrent irradiation of multiple targets at multiple discrete -ray beam energies. Our Au(,n) cross-section data for the production of Au in the ground state are in agreement with previous data measured using mono-energetic photon beams, but differ from data measured with a Bremsstrahlung beam above 14 MeV (see Figure 10 and Table 6). This observation is striking and deserves further investigation in future work. Furthermore, it can be seen that the evaluations based on TENDL and JENDL significantly underpredict the cross sections for energies above the giant dipole resonance (15MeV). This observation is likely to impact flux calibrations based on Au activation cross sections at these higher energies. In addition, experiments extending the measured cross sections further would provide further guidance to improve evaluations. Our data for the Au(,3n) reaction are given in Table 7 and plotted in Figure 11 in comparison to previous data and model calculations. As can be inferred from Figure 11, additional data are desirable to better evaluate the predictive power of the TENDL and JENDL approaches.
As discussed in Section I, the (,p) reaction is the primary pathway for producing high-specific-activity radioisotopes with electron LINACs. Our cross-section analysis can be extended to the remaining targets (TiO, Zn, and Os) to obtain (,p) reaction cross section. The main reactions of interest for the remaining targets are Ti(,p)Sc, Zn(,p)Cu, and Os(,p)Re. The data obtained for these reactions will be available in future publications.
VI.1 Incident photon scattering effect
The PHITS predictions of the scattering of the incident photons are presented in Figure 12. A few percentage points of the incident high-energy photons from the original beam are lost or scattered due to interactions with the atomic electrons in the target stack (Compton scattering), producing a lower-energy photon tail (see red histogram in Figure 12). Thus, the back Au target has an apparent 3 higher cross section than the first Au target in the beam energy range of 20-27 MeV. The effect increases the apparent cross section in the high-energy region by 2 mb, but has a very small effect in the giant resonance region where the cross sections are approximately 100 times larger, as shown in the simulations of Figure 12(b). Cross sections for the other targets in the stack will be corrected for this effect in future publications.
VI.2 Beam flux correction on the third segment
The photon beam at the outer radii (r 15 mm) for the third segment of each target was partially blocked by the copper apertures inside the vacuum chamber for the FEL aperture system of HIS. The beam flux on the third segment was found to be underestimated in the simulation. The correction factor of the beam flux on the third segment was determined to be 8. The cross sections for the third segments in Table 6 and 7 were adjusted to account for this effect.
VII Acknowledgments
This research was supported by the U.S. Department of Energy Isotope Program, managed by the Office of Science for Isotope RD and Production, under contract DE-AC02-06CH11357 (Argonne National Laboratory), and the Department of Energy, Office of Nuclear Physics under grant numbers DE-SC0018112, DE-SC0018325, DE-FG02-97ER41033 (TUNL) and DEFG02-97ER41041 (UNC). The authors would like to thank the HIS operators and staff for their help during the experiments. We gratefully acknowledge the computing resources provided on Bebop, a high-performance computing cluster operated by the Laboratory Computing Resource Center at Argonne National Laboratory, and the assistance of Dr. Sean Finch at TUNL in the analysis of targets used in this research.