### v2

# Photo-nuclear cross sections on 197Au, an update on the gold standard

Abstract
Cross sections for the 197Au reaction are broadly used in nuclear physics as a standard for normalizing photonuclear reaction cross-section data at photon beam energies above approximately 8 MeV. In this paper, we report cross-section measurements for the 197AuAug+m1 reaction at beam energies from 13 to 31 MeV. Our measurements provide the first cross-section data for this reaction at beam energies above 20 MeV, enabling the use of this reaction as a cross-section standard up to 30 MeV. Also, this work provides first cross-section measurements for the 197AuAum2 reaction. In addition, we measured cross-section data for the 197AuAu reaction, which can be used as a cross-section standard above about 25 MeV. These measurements were performed using a new target activation method that is based on the angle-energy correlation of the laser Compton-scattered photon beams at the High Intensity Gamma-ray Source (HIS). The technique enables measuring photonuclear reaction cross-sections at several discrete beam energies concurrently via a single irradiation on a stack of different targets. Measurements were carried out by irradiating a stack of concentric-ring targets consisting of Au, TiO2, Zn, Os, and Au (in order of the -ray beam direction). Our data for the 197Au(,n)196Aug+m1 reaction in the energy range of 13 to 20 MeV are in good agreement with existing ones measured using monoenergetic -ray beams, but differ from data acquired using a bremsstrahlung -ray beam. Also, above 18 MeV, our data for the 197Au(,n)196Aug+m1 and 197Au(,n)196Aum2 reactions differ significantly from the most recent TENDL and JENDL evaluations, suggesting a need to update these data libraries. The TENDL evaluation and existing data are consistent with our data for the 197Au(,3n) reaction, but differ significantly from the JENDL evaluation above 26 MeV.
I Introduction
Radioisotopes play important roles in numerous fields ranging from medical treatments to national security and basic research. Some examples include investigations of structures and reactions involving atomic nuclei, Mossbauer spectroscopy, radio-thermoelectric generation and other nuclear batteries, nuclear device detection, and the mitigation of nuclear proliferation [1]. The reports from “Workshop on the Nation’s Needs for Isotopes: Present and Future”, “Isotopes for the Nation’s Future, A long range plan”, and “Meeting Isotope Needs and Capturing Opportunities for the Future” identify multiple isotopes currently in short supply [1, 2, 3]. Among these are high specific activity beta emitters such as 47Sc, 67Cu, 77As, and 186Re. These radioisotopes are of interest to the community as they have ideal nuclear properties for medical applications. Their availability is limited as they are difficult to produce with high specific activity using common production methods.
Most radioisotopes are produced using either nuclear reactors or light-ion accelerators. High specific activity radioisotopes are very important in nearly all aspects of radiochemical work. Specific activity refers to the radioactivity of the produced radionuclide relative to the total mass of the element of interest present in the sample. Most radionuclides produced in nuclear reactors have low specific activity due to the inherently limited production pathways, e.g., dominated by (n,) reactions. In contrast, radionuclides produced at particle accelerator facilities via charged-particle induced reactions are generally characterized by high specific activity, e.g., (p,xn) reactions are typically used because the produced isotope is a different element from the target material, allowing chemistry techniques for separating materials. Also, photonuclear reactions, e.g., (,x), provide a promising process for producing high specific activity radioisotopes. The Nuclear Science Advisory Committee (NSAC) recently identified the use of photonuclear reactions as one of the most compelling opportunities for large-scale production of high specific activity radioisotopes using high-flux bremsstrahlung -ray beams generated with high-current linear accelerators (LINACs) [1]. The isotope production rate is directly proportional to the integral of the product of the -ray beam flux on the target and the photonuclear cross section. The end-point energy of the bremsstrahlung -ray beam spectrum is usually adjusted to optimize isotope production over the Giant Dipole Resonance (GDR) energy region. Accurate cross-section data for photonuclear reactions over this GDR energy region are needed for developing strategies for radioisotope production with high specific activity. A great deal of work has been done on measuring experimental cross-section databases for some reactions, e.g., (,n) reactions [4, 5, 6]. However, more measurements are needed. For example, the cross-section data for (,p) reactions (which lead to high specific activity radioisotopes) still require considerable work. There are many theoretical databases available which provide guidance on photo-production of radionuclides. Experimental data are needed to evaluate the accuracy of predictions of the leading reaction models used in nuclear-data libraries (TENDL, JENDL, ENDF, JEFF, CENDL, BROND, etc.) [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Examples of theoretically predicted cross sections which incorrectly estimate the production rate of radioisotopes are reported in the literature [17, 18, 19, 20]. Photonuclear cross sections are typically measured using broad energy spectrum bremsstrahlung -ray beams (either untagged or tagged) or quasi-monoenergetic -ray beams produced via laser Compton backscattering (LCB) or positron in-flight annihilation. For quasi-monoenergetic beams, LCB is the modern beam production method, see Table 1. Compared with a bremsstrahlung beam, which has a broadband energy intensity distribution, a LCB -ray beam is narrowly peaked around a specific energy. Precise measurement of cross sections can be performed using activation techniques with target irradiation by such LCB -ray beams. Independent of the type of -ray beam used, cross-section measurements require accurate determination of the beam flux. One approach is to establish a cross-section standard that can be used to determine the beam flux over a specified beam energy range. The cross sections for the 197Au(,xn) reactions are widely used as a standard for determining the photon beam flux in photonuclear reaction measurements. As a standard for cross-section normalization, errors in 197Au(,xn) cross-section data will impact a broad range of nuclear physics research and applications. For example, gold monitor foils are routinely used to normalize bremsstrahlung photon beams in cross-section measurements [21, 22, 23]. The current status of 197Au(,x) cross-section data is as follows. The 197Au(,n)196Au cross section data from different experiments are in good agreement in the energy range of 8-14 MeV, which is below the threshold for multiple neutron emission (see Table 1). However, there are significant discrepancies between data [24, 25] at photon energies between 14-18 MeV. Also, there are no cross-section data for the 197Au(,n) reaction above 20 MeV and extrapolation of data evaluations by TENDL and JENDL differ substantially.
In this paper, we describe our newly developed experimental techniques for measuring photonuclear cross sections at discrete energies using a LCB -ray beam and its application to measuring cross sections for photon-induced reactions on 197Au. The photon energies in this work were chosen to be above the threshold for reactions leading to high specific activity radionuclides via the (,p) reaction for use in medical applications such as 48Ti(,p)47Sc and 68Zn(,p)67Cu at around 20 MeV and higher. The technique is based on the well-defined radial energy dependence of a LCB -ray beam which enables measurement of the corresponding spatially dependent cross section. The central portion of the -ray beam has the highest energy and flux density, with the beam energy and flux density decreasing radially outward from the center in a predictable manner [26]. Activation of concentric ring targets by an uncollimated beam will provide experimental activation data to be used to determine cross sections at the average energy of the beam incident on each ring. This method was developed and used for the first time at the Triangle Universities Nuclear Laboratory’s (TUNL) High-Intensity Gamma-Ray Source (HIS). The HIS facility is described by Weller et al. in Ref. [27]. A goal of the present work is to measure the exclusive cross-section excitation functions of the 197Au(,n) and (,3n) reactions for use as cross-section standards. Table 1 provides a summary of the world data for the 197Au(,xn) reactions by beam type. An interesting feature of the existing data is that there is apparently a strong correlation between the beam type (quasi-monoenergetic or bremsstrahlung) and the experimental technique. All experiments in Table 1 using a quasi-monoenergetic -ray beam measured the cross section with direct neutron counting. The measurements using a bremsstrahlung -ray beam determined the cross section via activation techniques. Our measurements are the first to be done with a monoenergetic -ray beam while using activation techniques. In Section V, our data are compared to the results of the previous measurements listed in Table 1. The results of this work for the Ti, Zn, and Os targets measured concurrently with Au will be presented in a subsequent paper. Here, our multiple target multi-energy technique is described in section II, the HIS LCB beam is presented in section III, the experiment is discussed in section IV, section V describes the cross-section analysis, and the results are presented and discussed in section VI.
| Reaction | Reference | Year | range (MeV) | Points | Lab | Beam type | Technique |
| (,n) | C. Plaisir et al. [24] | 2012 | 10-20 | 101 | CEA/DAM ELSA, France | Bremsstrahlung | activation counting |
| (,n) | K. Vogt et al. [28] | 2002 | 8.08-9.90 | 9 | TU-Darmstadt, Germany | Bremsstrahlung | activation counting |
| (,n) | B.L. Berman et al. [29] | 1987 | 12.1-16.9 | 11 | LLNL, USA | annihilation | neutron counting |
| (,n) | A. Veyssiere et al. [30] | 1970 | 8.08-19.8 | 43 | CEA Paris-Saclay, France | annihilation | neutron counting |
| (,n) | S.C. Fultz et al. [25] | 1962 | 8.7-17.9 | 30 | AIST, Tsukuba, Japan | annihilation | neutron counting |
| (,n) | O. Itoh et al. [31] | 2011 | 8.08-13.14 | 30 | AIST, Tsukuba, Japan | LCB source | neutron counting |
| (,n) | F. Kitatani et al. [32] | 2011 | 9.3-12.4 | 12 | AIST, Tsukuba, Japan | LCB source | neutron counting |
| (,n) | F. Kitatani et al. [33] | 2010 | 10.0-14.6 | 8 | AIST, Tsukuba, Japan | LCB source | neutron counting |
| (,n) | K.Y. Hara et al. [34] | 2007 | 8.23-12.3 | 24 | AIST, Tsukuba, Japan | LCB source | neutron counting |
| (,2n) | B.L. Berman et al. [29] | 1987 | 14.0-16.9 | 7 | LLNL, USA | annihilation | neutron counting |
| (,2n) | A. Veyssiere et al. [30] | 1970 | 13.5-27.1 | 40 | CEA Paris-Saclay, France | annihilation | neutron counting |
| (,2n) | S.C. Fultz et al. [25] | 1962 | 14.3-24.7 | 29 | LLNL, USA | annihilation | neutron counting |
| (,3n) | A. Veyssiere et al. [30] | 1970 | 22.8-27.1 | 8 | CEA Paris-Saclay, France | annihilation | neutron counting |
II Method
This section provides details about the method applied in this work to measure the exclusive (,n) and (,3n) cross sections via the activation method. The excitation functions reported here cover the -energy range from 13 to 31 MeV. The photonuclear cross sections over the range of the Giant Dipole Resonance (GDR) vary by orders of magnitude. The goal of this work is to demonstrate our activation techniques by applying it in measuring the excitation functions of the 197Au(,xn) reaction cross sections. with an energy resolution of 3-10 at each energy. At HIS, the beam energy spread depends on the diameter of the beam defining collimation. In a typical setup with an energy spread (FWHM) of 5 (FWHM), the flux transmitted through the collimator is about 7.5 of the total flux from the Compton-scattered laser beam. Our method of using three radial target segments enables concurrent measurement of the cross section at three central beam energies. The radial energy and flux distribution of the HIS beam was verified experimentally by blocking portions of the beam with tungsten plugs to mimic the concentric-rings target design. Simulations were performed and compared to the experimental results of these calibrations to determine the accuracy of the beam simulations.
II.1 Target stack
In the activation method, the photon beam induces radioactivity via photon-induced reactions, such as (,n) and (,p), in targets. Reaction yields are determined by gamma counting the targets after activation. The target foils used in these experiments are made of separable concentric rings that exploit the energy spread of the HIS beam and enable the activation of multiple targets across multiple energies in a single beam exposure. A concept of a target stack is illustrated in Figure 1. The relatively low and well-characterized attenuation of high-energy photons by the target materials provides the ability to irradiate multiple targets in a single irradiation. The total thickness of the stack of targets used in these experiments was 2.8-3.3 g/cm2, which resulted in a 13 total attenuation of the photon beam at 14 MeV. The entire target may then be exposed to the beam and each individual concentric ring will interact with a narrow range of beam energies and a calculable fraction of the overall photon flux. The individual rings are analyzed with high-purity germanium detectors (HPGe) to determine activation yields. Thus, using this technique, the cross sections at three energies can be measured in a single irradiation. With the stack of five targets and three segments per target, yields for 15 target segments are obtained for each 12-hour run.
II.2 Activation
Using the activation method (see details in Appendix A), the cross sections on the 197Au targets were obtained. The combined cross section for ground and isomeric state is given by,
with,
The cross section for producing 196Au in the 2nd isomer state is,
The cross section for producing 194Au is determined in the same way as Equation 2,
III quasi-monoenergetic Photon beam
The photon beams are produced by Compton scattering of photons from relativistic electrons circulating in a storage ring at HIS. In the case of relativistic electrons ( 1) scattering at a small angle (1), the scattered photon energy after collision is given by,
where Eg is the scattered photon energy, and Ep and Ee are the energies of the incident photon and electron. The Lorentz factor of the electron is denoted as . The angle is expressed in terms of the radius of the target (r) and the distance from the collision point to the target position, r/L.
As stated above, for each beam irradiation, reaction cross sections were measured for each target at three average beam energies, one for each target segment. The cross sections were determined from the measured activation of each target segment using Equations 14 – 16 (see Appendix A). The parameter in the equations is the -ray beam flux incident on each target segment, i.e., , where is the total beam flux in the direction of the target and is the fraction of incident on ring i. The uncertainty in , , is the dominant systematic uncertainty in the cross-section measurements. There are two systematic uncertainties contributing to : (1) the uncertainty in the measurement of , and (2) the uncertainty in the calculated value of . The measurements are described in section IV.2. The determination of for each segment and its uncertainty is described in this section. We used the model of Sun and Wu [26] to simulate the Compton-scattered photons passing through the free-electron laser (FEL) aperture systems to obtain the beam intensity profile (i.e., , the beam intensity as a function of the radial distance from the beam axis) and the photon energy spectrum (i.e., , the beam intensity as a function of the photon energy . The calculated ) and from the simulations have been benchmarked with other codes such as CAIN by Yokoya [35] and using experimental measurements at HIS [26]. An example of the Monte-Carlo simulations for our experiment is given in Figure 2. For this simulation, the -ray beam at HIS is produced with 660 MeV and a free electron laser photon wave length nm. The radial extent of each ring is given in Figure 2. The energies of peak intensity of the scattered -ray beam at the target location (L 53 m) incident on the three segments are 18, 17 and 16 MeV, respectively. The beam intensity distributions incident on each ring, and , are plotted in Figure 2 a) and b), respectively. For the beam energy range of 13 to 31 MeV covered in our measurements, the typical percentile energy spreads (1) of the photon beam incident on the , and target rings were 1.3-2.7, 2.2-4.3 and 3.2-5.7 of the central beam energy, respectively. The and distributions were computed from the simulations and used to calculate the values of and , the photon energy of the peak intensity of the distribution for each ring. We performed benchmark measurements to assess the systematic uncertainties in the calculations of and using the simulated data. These measurements were performed using rays produced under conditions identical to those used in the simulations, MeV and nm. Special beam collimator arrangements were used to select sections of the beam cross section like the rings in the targets. Each collimator configuration consisted of a lead collimator with an open circular aperture, or the same lead collimator with a circular tungsten rod insert. The lead collimator and tungsten rods were 6 inches (15.2 cm) in length. The axes of the collimator aperture and the cylindrical rod were centered on the beam axis. At the central beam energy of about 25 MeV, the -ray attenuation through the body of the lead collimators and tungsten rods were about and , respectively. The energy spectrum of the beam was measured using a large cylindrical NaI detector (10” dia. 14” thick.), referred to as “Molly”. The efficiency of Molly is greater than 98 at the energies of these measurements. For the beam energy profile measurements, Molly was positioned with its axis centered on the -ray beam axis, and copper attenuators were inserted in the beam path to reduce the beam flux incident on Molly to less than about , where the pulse-pileup rate is negligible. Measurements of were performed for the first four collimator configurations itemized in Table 2 a): N1 = 0.0 to 6.35 mm; N2 = 0.0 to 15.9 mm; N3 = 6.35 to 15.9 mm; and N4 = 11.1 to 15.9 mm. These measurements were carried out using two collimators with aperture diameters of 0.5” (12.7 mm) and 1.25” (31.8 mm) and two tungsten rods with diameters 0.5” (12.7 mm) and 0.875” (22.2 mm). In addition, the difference in the measured for two collimator configurations was determined by comparing measured data to simulations for radial ranges similar to the target rings, e.g., see configuration 5 in Table 2. The values for obtained from the simulated distributions differ from those determined using the measured distributions by about 2, see the results for collimator configurations 1–4 in Table 2. This uncertainty is significantly smaller than the error in our determination of the absolute beam flux, see section IV.2. As shown in Table 2, the mean beam energy computed using the simulated data is in good agreement with the measured mean beam energy in each segment. Examples of measured distributions for several collimator configurations can be found in Figure 3.
| Coll. | Coll. Dia. | Rod Dia. | Meas. | Sim. | Meas. | Sim. | Note | ||
| Config. | mm | mm | mm | mm | MeV | MeV | |||
| 1 | 12.7 | none | 0.0 | 6.35 | 18 | 20 | 27.1 | 27.1 | |
| 2 | 31.8 | none | 0.0 | 15.9 | 100 | 100 | 25.2 | 25.6 | (a) |
| 3 | 31.8 | 12.7 | 6.35 | 15.9 | 81 | 80 | 24.9 | 25.6 | |
| 4 | 31.8 | 22.2 | 11.1 | 15.9 | 45 | 45 | 24.1 | 24.5 | |
| 5 | 31.8 | 12.7/22.2 | 6.35 | 11.1 | 37 | 35 | 25.8 | 26.1 | (b) |
IV Experiment
Targets were activated at seven different central photon beam energies (see Table 3). The gamma energy at the outermost ring typically was close to the central energy of the next lower energy irradiation, thus providing an internal consistency check for these measurements. Using targets with three concentric segments, the cross-section measurements for the reactions 197Au(,n)196Au and 197Au(,3n)194Au were determined for 21 different energies. The target stack was irradiated for 8-12 hours by a photon beam with fluxes ranging between 108-109 /s, limited by a copper pre-collimator. The boundaries of the photon beam of 30-40 mm in diameter were shaped by the copper apertures inside the HIS aperture systems and by the copper pre-collimator. Irradiation times, beam intensities and uncertainties are summarized in Table 3.
| Irradiation time | Intensity | Unc. | |
| MeV | hr | /s | |
| 14 | 11.49 | 8.01 | 8.2 |
| 16 | 10.72 | 8.80 | 8.2 |
| 18 | 8.17 | 9.27 | 8.2 |
| 22 | 10.73 | 2.48 | 8.2 |
| 24 | 11.20 | 2.70 | 8.2 |
| 27 | 11.73 | 2.51 | 8.2 |
| 31 | 9.12 | 2.06 | 8.2 |
IV.1 Target stack
The target stack was composed of layers of Au, TiO2, Zn, Os and Au as illustrated in Figure 4 and listed in Table 4. The target segments were mounted on 3D-printed acrylic with cavities made for receiving powdered target material. The cavities were filled with a known mass of the target material and then sealed by adding and curing 3D printing gel on top of this target material. The effects of photon attenuation and scattering by the low-acrylic target mounts were negligible. The thickness of the targets varied across the numerous ones made for this work (see Table 4). The front and back Au targets were arranged in this manner in anticipation of using these to normalize the beam flux and calibrate the beam energy throughout the stack. Placed front to back in the stack, relative to the incident beam, and between the front and back Au targets, were TiO2, Zn and Os ones. The beam flux and scattered gamma contribution to the cross section were calculated. There were 3 sets of targets to enable reuse after sufficient decay of the activations in subsequent runs separated by several months.
| 14, 27 | 16, 24 | 18, 22, 31 | |||
| thick. | |||||
| tgt. | seg. | (g/cm2) | (g/cm2) | (g/cm2) | |
| 1 | 197Au | 1 | 0.681 | 0.713 | 0.763 |
| 2 | 0.931 | 0.933 | 0.655 | ||
| 3 | 1.033 | 0.935 | 0.869 | ||
| 2 | TiO2 | 1 | 0.228 | 0.241 | 0.235 |
| 2 | 0.205 | 0.194 | 0.163 | ||
| 3 | 0.205 | 0.220 | 0.221 | ||
| 3 | Zn | 1 | 0.501 | 0.350 | 0.308 |
| 2 | 0.392 | 0.292 | 0.363 | ||
| 3 | 0.351 | 0.339 | 0.406 | ||
| 4 | Os | 1 | 0.304 | 0.293 | 0.358 |
| 2 | 0.257 | 0.242 | 0.220 | ||
| 3 | 0.614 | 0.553 | 0.362 | ||
| 5 | 197Au | 1 | 1.209 | 1.299 | 1.115 |
| 2 | 0.834 | 1.195 | 1.049 | ||
| 3 | 1.304 | 1.395 | 1.471 |
IV.2 Photon beam flux
A significant portion of the overall systematic uncertainty in the cross sections measured in this work is due to the errors in measuring the absolute -ray beam flux. A schematic diagram of the system used for this purpose during target irradiation is given in Figure 5. The beam flux was measured using three detectors during target irradiation: 1) a shielded plastic scintillator paddle which detects rays and recoil electrons from Compton scattering of the beam passing through the mirror of the optical cavity (labeled MP in Figure 5), 2) a thin plastic scintillator centered in the beam path just after the collimator (labeled SP in Figure 5), and 3) a silicon detector which measures charged particles emitted from photon-induced reactions in a thin CD2 foil inside an evacuated scattering chamber (labeled CP in Figure 5). The MP detector monitors a portion of the full -ray beam intensity passing through the cavity mirror. The absolute detection efficiency of MP for measuring the beam passing through the collimator is determined using Molly (see Figure 5). The Molly detector is described earlier. Photon interactions in Molly are modeled with GEANT-4. According to these GEANT-4 simulations, the full-energy peak efficiency of Molly is greater than 98 at energies above 10 MeV. Molly can be positioned in the beam path, and a system of high-precision copper absorbers can be inserted into the beam to attenuate the incident flux on Molly to keep the data acquisition system deadtime lower than 40 and the pulse pileup rate of the detector signal below 10. The efficiencies of the other flux monitors, SP and CP, are calibrated relative to MP using the full beam intensity. In this arrangement, the ratios of the rates in SP and CP to that in MP are computed to monitor the stability of the fraction of the total -ray flux that is transmitted through the collimator. Over the duration of each run, the efficiency of MP is determined using the following relationships,
where . The symbol is the photon attenuation coefficient (cm2/g) averaged over the beam intensity profile, is the density of the copper attenuators (8.96 g/cm3), and x is the length in cm of the copper attenuator inserted in the beam path to reduce the flux incident on Molly, while is the net counting rate in the MP paddle. The symbol denotes the -ray energy where the distribution has a maximum value, see Figure 6. The value of was calculated using the photon attenuation coefficients as a function of energy, (E), obtained from NIST [36, 37], and the beam intensity profile simulated as described in [27]. An example of and , the simulated unattenuated and attenuated beam intensity profiles, used to calculate for the beam energy distribution with the peak beam energy, = 18.0 MeV, is plotted in Figure 6. The value is the -ray energy, E, where the distribution has a maximum value. The distribution is normalized such that the integration over the distribution is equal to 1.0. The main systematic uncertainty in determining in the above equation is due to the error in the value of the attenuation . The relative uncertainty in the measured is less than 1 and is not included in the derivation of . The relative uncertainty in is,
The calculated value was benchmarked at several beam energies with in-situ measurements. The average variance between the calculated and measured values was 0.7. We use this variance as an estimate of the uncertainty in the calculated value. In turn, the uncertainty in was used to compute the average error in the calculated intensity. For these measurements, the relative systematic uncertainty in the -ray beam flux is = (0.007)(8.96 g/cm3)(0.033 cm2/g)(39 cm) = 0.081 = 8.1. The relative uncertainty in transfers to the uncertainty in measured MP efficiency, see equation 5. The statistical uncertainty in is less than 1. The efficiency of MP as a function of is displayed in Figure 7. The error bars are the combined statistical and systematic uncertainties.
IV.3 The activated target -ray decay spectra
After activation, -ray spectra were measured with 12 High Purity Germanium (HPGe) detectors, each with a co-axial cylindrical crystal positioned with its front face 5 cm from the activated target. The measurements started approximately 30 minutes after the end of bombardment and lasted long enough to obtain sufficient counts for the detection of the second isomeric state of 196Au with the half-life of 9.6 hours. Typical HPGe detector efficiency shown in Figure 8 was calibrated with a mixed calibration source, and the detector efficiencies for the gold isotopes were determined by the formula [39].
where x and y are ln and ln/1000), respectively, Eγ is the -ray energy from selected nuclei. The absolute detection efficiency at 333 keV for the 196Au transition is 1-3, and the corresponding uncertainty, is 2. During the measurements, the dead time was typically less than 2. The spectra at the photon beam energies of 18 and 27 MeV are presented in Figure 8. The -ray peaks of 147.81, 188.27 keV (second isomeric state of 196Au), 333.03, 355.73 keV (ground state of 196Au) and 293.55, 328.47 keV (194Au) are visible in Figure 8.
V Cross-section analysis
The photonuclear cross sections were determined using equations 14 and 15. In addition, the detector efficiency, detector dead time, branching ratio, beam attenuation and self-absorption effects are taken into account. With these corrections, the cross sections can be expressed as,
where and are the cross sections for the ground + 1st isomeric state and 2nd isomeric state, respectively. stands for the detected counts, and are the beam flux and the detector efficiency, respectively. The total correction factor above is expressed as = . The , , , and symbols are corrections for the detector efficiency, detector dead time during the measurements, -ray branching ratio, -ray beam attenuation and self-absorption by passing through the targets, respectively. Correction of the primary photon beam flux due to attenuation, and of the emitted -ray yield from Au isotopes due to self-absorption are taken into account by,
where and x0 are attenuation coefficients for the beam and emitted -ray and the total length of the target, respectively. The statistical error was calculated by the RADWARE software [39] and three systematic errors for the -ray branching ratio, detector efficiency and beam flux were determined. Therefore, the total experimental error can be obtained by,
where , , and are uncertainties from the number of nuclei measured by counts, photon beam flux, HPGe detector efficiency and the branching ratio of the selected nuclei, respectively. The uncertainties of the branching ratios are given in Table 5.
After activation by the photon beam, the segments in the target stacks were separated and counted individually with HPGe detectors. The perceptible Au radioisotopes produced via 197Au(,xn) were 196Au and 194Au. The 196Au nucleus was observed to produce two isomeric states (m1= 5+, m2= 12-) and the ground state (2-). The simplified decay schemes for the 196Au and 194Au nuclei are illustrated in Figure 9. The half-life of the , and ground states of 196Au are 8.1 s, 9.6 h and 6.17 d, respectively. The m2 and ground states are measurable and distinguishable by counting the rays from the gold targets within a few hours from the end of each irradiation period. The m1 isomeric state, with its short half-life of only 8.1 s, has its yield included in the decay of the ground state. The 147.81- and 188.24-keV -ray lines shown in Figure 9 and the peaks in Figure 8 are used to determine the yield of the m2 isomer, and the 333.03-, 328.47-keV -ray transitions following emission from the ground state of 196Au are used to calculate the cross section for the sum of the ground state and m1 isomer, respectively. The ground state (1-) of 194Au has a 38.02 h half-life and decays by emission to 194Pt. The 293.55 and 328.47 keV -ray peaks were used to determine the cross section of the 197Au(,3n)194Au reaction. The decay information for the 196Au and 194Au nuclei, along with uncertainties, is listed in Table 5.
| nucl. | decay mode | half-life | energy | branching ratio |
| (keV) | () | |||
| 196Aug | 93.0 | 6.1669 (6) d | 333.03 | 22.9 (9) |
| 355.73 | 87 (3) | |||
| 196Aum2 | IT 100 | 9.6 (1) h | 147.81 | 43.5 (15) |
| 188.27 | 30.0 (15) | |||
| 194Aug | 100 | 38.02 (10) h | 293.549 | 10.9 (3) |
| 328.470 | 62.8 (16) |
V.1 197Au(,n)196Au
The cross sections of the combined ground and first isomeric states and second isomeric state for the photon-induced reaction on 197Au are obtained with the emitted rays of 333.03, 355.73 keV and 147.81, 188.27 keV, respectively. The excitation functions for the 197Au(,n)196Au and 197Au(,n)196Au reactions are presented in Figure 10 together with the previous experimental data [24, 25, 34, 33, 31, 32, 30, 29, 28] and theoretical calculations from TENDL 2021 and JENDL 4.0. Here, the cross sections in Figure 10 were determined by means of weighted mean and error. The cross sections measured in this work with statistical and total uncertainties are given in Table 6.
| nucl. | 196Au | 196Au | |||||||||||||||
| peak | 333.03 | keV | 355.73 | keV | 147.81 | keV | 188.27 | keV | |||||||||
| stat. | tot. | stat. | tot. | stat. | tot. | stat. | tot. | ||||||||||
| MeV | MeV | mb | mb | mb | mb | mb | mb | ||||||||||
| 12.95 | 0.41 | 518.26 | 0.20 | 9.31 | 512.33 | 0.10 | 9.12 | 514.96 | 8.72 | 0.16 | 63.6 | 64.3 | 0.16 | 50.4 | 51.4 | 0.16 | 40.27 |
| 13.48 | 0.30 | 545.27 | 0.20 | 9.31 | 541.72 | 0.10 | 9.12 | 543.30 | 8.72 | 0.24 | 2.3 | 9.4 | 0.25 | 38.4 | 39.7 | 0.24 | 9.39 |
| 13.96 | 0.18 | 533.03 | 0.20 | 9.31 | 530.82 | 0.10 | 9.12 | 531.81 | 8.72 | 0.28 | 16.9 | 19.2 | 0.27 | 9.2 | 13.5 | 0.27 | 10.85 |
| 14.58 | 0.47 | 500.85 | 0.20 | 9.31 | 496.38 | 0.10 | 9.12 | 498.37 | 8.72 | 0.37 | 5.3 | 10.5 | 0.35 | 9.9 | 14.0 | 0.36 | 9.73 |
| 15.30 | 0.43 | 370.52 | 0.20 | 9.31 | 367.09 | 0.10 | 9.12 | 368.62 | 8.72 | 0.41 | 8.3 | 12.3 | 0.44 | 5.8 | 11.4 | 0.43 | 8.67 |
| 15.94 | 0.23 | 241.49 | 0.30 | 9.31 | 240.44 | 0.20 | 9.12 | 240.91 | 8.72 | 0.39 | 6.1 | 11.0 | 0.42 | 6.4 | 11.7 | 0.41 | 8.91 |
| 16.18 | 0.64 | 193.69 | 0.40 | 9.32 | 192.63 | 0.20 | 9.12 | 193.10 | 8.72 | 0.30 | 7.7 | 11.9 | 0.33 | 6.6 | 11.8 | 0.33 | 9.06 |
| 17.10 | 0.49 | 93.08 | 0.50 | 9.33 | 92.25 | 0.30 | 9.12 | 92.61 | 8.72 | 0.20 | 5.8 | 10.8 | 0.25 | 6.4 | 11.7 | 0.23 | 8.92 |
| 17.91 | 0.24 | 54.69 | 0.60 | 9.33 | 55.28 | 0.30 | 9.12 | 55.01 | 8.72 | 0.20 | 6.1 | 11.0 | 0.19 | 5.7 | 11.3 | 0.19 | 8.59 |
| 19.85 | 0.82 | 20.26 | 2.00 | 9.53 | 21.09 | 1.10 | 9.18 | 20.73 | 8.78 | 0.12 | 14.0 | 16.7 | 0.13 | 14.1 | 17.2 | 0.12 | 12.78 |
| 20.90 | 0.69 | 15.27 | 4.10 | 10.16 | 14.69 | 0.70 | 9.14 | 14.86 | 8.86 | 0.12 | 9.3 | 13.0 | 0.11 | 17.4 | 20.0 | 0.12 | 11.90 |
| 21.83 | 0.48 | 13.62 | 3.50 | 9.70 | 12.64 | 1.20 | 9.26 | 13.00 | 8.84 | 0.10 | 7.0 | 14.3 | 0.10 | 11.3 | 16.9 | 0.10 | 11.92 |
| 21.72 | 0.99 | 12.69 | 2.70 | 9.94 | 12.23 | 1.60 | 9.20 | 12.38 | 8.86 | 0.10 | 11.0 | 11.5 | 0.09 | 13.7 | 15.0 | 0.09 | 10.34 |
| 22.99 | 0.83 | 11.25 | 2.70 | 9.71 | 10.40 | 4.50 | 9.43 | 10.74 | 8.90 | 0.09 | 16.0 | 18.4 | 0.09 | 7.1 | 12.1 | 0.09 | 9.50 |
| 24.12 | 0.57 | 8.28 | 4.20 | 11.78 | 8.52 | 1.40 | 9.53 | 8.46 | 9.22 | 0.09 | 43.2 | 15.4 | 0.08 | 23.6 | 10.8 | 0.08 | 10.26 |
| 23.43 | 1.13 | 9.28 | 7.20 | 10.21 | 9.31 | 2.80 | 9.22 | 9.30 | 8.90 | 0.09 | 12.4 | 44.2 | 0.08 | 4.5 | 25.6 | 0.08 | 21.96 |
| 24.91 | 0.95 | 7.96 | 3.30 | 9.89 | 8.16 | 1.40 | 9.23 | 8.08 | 8.86 | 0.09 | 43.1 | 44.1 | 0.08 | 15.2 | 18.1 | 0.08 | 15.94 |
| 26.23 | 0.66 | 5.65 | 9.00 | 12.95 | 6.23 | 7.40 | 11.77 | 5.96 | 10.44 | 0.08 | 32.6 | 33.8 | 0.07 | 16.1 | 18.8 | 0.07 | 16.10 |
| 27.24 | 1.51 | 5.22 | 2.60 | 9.67 | 4.28 | 0.50 | 9.13 | 4.56 | 8.79 | 0.08 | 19.8 | 21.8 | 0.06 | 29.7 | 31.3 | 0.07 | 18.55 |
| 29.25 | 1.26 | 3.83 | 3.50 | 9.95 | 3.89 | 1.70 | 9.28 | 3.87 | 8.89 | 0.07 | 41.5 | 42.5 | 0.07 | 20.1 | 22.4 | 0.07 | 19.48 |
| 31.05 | 0.84 | 3.21 | 4.40 | 10.31 | 3.02 | 1.70 | 9.28 | 3.08 | 8.94 | 0.05 | 18.3 | 20.4 | 0.05 | 17.3 | 19.9 | 0.05 | 14.87 |
V.2 197Au(,3n)194Au
The cross sections for the 197Au(,3n)194Au reaction are obtained from the emitted rays of 293.55 and 328.47 keV. The weighted mean cross sections are presented in Figure 11, together with the previous experimental data [30] and theoretical calculations from TENDL 2021 and JENDL 4.0. The cross sections measured in this work with statistical and total uncertainties are summarized in Table 7.
| nucl. | 194Au | ||||||||
| peak | 293.55 | keV | 328.46 | keV | |||||
| stat. | tot. | stat. | tot. | ||||||
| MeV | MeV | mb | mb | mb | |||||
| 22.99 | 0.83 | - | - | - | 0.12 | 5.31 | 11.5 | 0.12 | 11.5 |
| 23.43 | 1.13 | 0.86 | 6.3 | 10.7 | 0.90 | 2.13 | 8.8 | 0.89 | 8.7 |
| 24.12 | 0.57 | 1.83 | 7.6 | 9.0 | 2.00 | 1.22 | 8.6 | 1.93 | 8.5 |
| 24.91 | 0.95 | 7.63 | 2.2 | 8.8 | 8.02 | 0.92 | 8.6 | 7.86 | 8.4 |
| 26.23 | 0.66 | 16.18 | 0.9 | 8.6 | 17.02 | 1.91 | 8.8 | 16.51 | 8.4 |
| 27.24 | 1.51 | 16.24 | 1.7 | 8.7 | 15.78 | 0.44 | 8.6 | 15.96 | 8.4 |
| 29.25 | 1.26 | 21.25 | 1.2 | 8.6 | 24.54 | 0.39 | 8.6 | 22.86 | 8.4 |
| 31.05 | 0.84 | 20.41 | 1.3 | 8.7 | 21.70 | 0.41 | 8.6 | 21.10 | 8.4 |
VI Results and Discussion
This work demonstrates the successful development of a novel and highly efficient method for concurrently measuring photonuclear reaction cross sections at multiple discrete photon energies for various materials using a single irradiation. As a proof-of-concept, the excitation function of the 196Au(,n)196Au and 196Au(,3n)194Au reactions were measured with this novel method utilizing target stacks consisting of concentric ring segments bombarded with an uncollimated beam with well-defined radial energy and flux distributions. The measurements were performed at 21 discrete beam energies over the incident photon beam energy range of 13-31 MeV (see Table 6). For the first time, the cross sections of the combined ground and first isomeric states at beam energies above 20 MeV and the second isomeric state for the 197Au(,n)196Au reactions, were determined via the activation method which yielded exclusive cross sections which are unaffected by multiple neutron emission. The method developed here based on radial target segmentation and target stacking makes efficient use of beam time by concurrent irradiation of multiple targets at multiple discrete -ray beam energies. Our 197Au(,n) cross-section data for the production of 196Au in the ground state are in agreement with previous data measured using mono-energetic photon beams, but differ from data measured with a bremsstrahlung beam above 14 MeV (see Figure 10 and Table 6). Our data, which were measured with a mono-energetic -ray beam using activation techniques, are in good agreement with data measured with a similar beam but using neutron counting techniques. This result suggests that the standing discrepancies between data are most likely associated with beam type and not experimental techniques, i.e., neutron counting versus target activation. Furthermore, the TENDL and JENDL evaluations significantly underpredict the cross sections at energies above 18 MeV, suggesting a need to update these evaluations. Our data for the 197Au(,3n) reaction are given in Table 7 and plotted in Figure 11 in comparison with previous data and model calculations. As can be inferred from Figure 11, additional data at energies above 30 MeV are desirable to better estimate the predictive powers of the TENDL and JENDL evaluations. As discussed in Section I, the (,p) reaction is the primary pathway for producing high specific-activity radioisotopes using bremsstrahlung rays generated by electron LINACs. Our cross-section analysis can be extended to the remaining targets (TiO2, Zn, and Os) to obtain (,p) reaction cross sections. The main reactions of interest for the remaining targets are 48Ti(,p)47Sc, 68Zn(,p)67Cu, and 189Os(,p)188Re. The data obtained for these reactions will be reported in future publications.
VI.1 Incident photon scattering effect
Having accurate values for the beam flux at each segment of each target is essential in determining the corresponding reaction cross sections. The fractional photon beam fluxes at each target segment were obtained from simulations, which were benchmarked with measurements as described above. With these incident beam fluxes on the segments of the front target, the beam intensity at each target was calculated using the beam transport simulation performed by the PHITS Monte-Carlo code [41]. The geometry illustrated in Figure 4 was implemented in the simulations. The targets in the stack were Au, TiO2, Zn, Os and Au. Successive targets were separated by 0.45 cm. The detailed information about the segments of the actual target stack can be found in Section IV. A total of 15 segments in the target stack (3 segments per target) were exposed to the photon beam at once. The beam flux was simulated for each irradiation (See Table 8). In the energy range of our measurements, Compton scattering and pair production dominate the beam attenuation and scattering as it passes through the stack. The attenuation of the gamma beam is simulated in detail using the PHITS code with the overall attenuation predicted to be 10. The central beam energy, one-sigma radial energy spread, and fractional beam flux on each segment of the stack are given in Table 8. The overall simulation is validated using the data from the gold target at the back of the stack when compared with the data from the one in the front. Predictions of distortions to the photon energy spectrum due to the photons scattering as they pass through the target stack are made using PHITS. Some results from the PHITS calculations are found in Figure 12. In these calculations is about 25 MeV, as can be inferred from the shaded peak in Figure 12 a). As the photon bunch traverses the target stack, the energy of a small fraction of the photons is degraded due to Compton scattering from atomic electrons. This effect creates the low-energy tail on the photon energy distribution (red curve) displayed in Figure 12 a). Because this low-energy tail extends through the GDR region, the cross section measured with the back Au target is higher than that measured using the front one in the beam energy range of 20-27 MeV. This effect increases the apparent cross section in the high-energy region by 2 mb, but has a very small effect in the giant resonance region where the cross sections are approximately 100 times larger, as shown by the simulations plotted in Figure 12 b). Cross sections for the other targets in the stack will be corrected for this effect in future publications.
VI.2 Beam flux correction on the third segment
The photon beam at the outer radii (r 15 mm) for the third segment of each target was partially blocked by the copper apertures inside the vacuum chamber for the FEL aperture system of HIS. The beam flux on the third segment was found to be underestimated in the simulation. The correction factor of the beam flux on the third segment was determined to be 8. The cross sections for the third segments in Table 6 and 7 were adjusted to account for this effect.
VII Conclusions
We report cross-section measurements for the 197Au(,n)196Aug+m1, 197Au(,n)196Aum2, and 197Au(,3n)194Au reactions at -ray beam energies ranging from 13 to 31 MeV. The measurements were performed at HIS demonstrating the enhanced efficacy of a new sample activation method based on irradiating targets consisting of concentric-ring segments relative to using solid targets. In this proof-of-concept experiment, data at three discrete -ray energies were collected for each target irradiation. Our results represent the first cross-section data for the 197Au(,n)196Aug+m1 reaction above 20 MeV. As such, they enable the first data-based assessment of the predictions of the TENDL [7] and JENDL [8, 9, 10, 11] evaluated nuclear data libraries above 20 MeV for this reaction and provide new data for updating the reaction models used in the evaluations. Also, our data for the 197Au(,n)196Aum2 reaction, leaving 196Au in the = metastable state ( = 9.6 h), are the first, and thereby provide constraints for the modeling of this reaction in TENDL. These new measurements have significant implications on the use of the 197Au(,n)196Aug+m1 reaction as a cross-section standard. Our data for the 197Au(,n)196Aug+m1 reaction are in good agreement with TENDL predictions and existing data that were measured using a monoenergetic -ray beam and neutron counting techniques. However, they differ by as much as 60 from data measured with a bremsstrahlung beam at energies between 13 and 16 MeV. This observation suggests that, in the energy range of 13 – 16 MeV, preference should be given to using cross sections for the 197Au(,n)196Aug+m1 reaction that are consistent with data measured using monoenergetic -ray beams.
VIII Acknowledgments
This research was supported by the U.S. Department of Energy Isotope Program, managed by the Office of Science for Isotope RD and Production, under contract DE-AC02-06CH11357 (Argonne National Laboratory), and the Department of Energy, Office of Nuclear Physics under grant numbers DE-SC0018112, DE-SC0018325, DE-FG02-97ER41033 (TUNL) and DEFG02-97ER41041 (UNC). The authors would like to thank the HIS operators and staff for their help during the experiments. We gratefully acknowledge the computing resources provided on Bebop, a high-performance computing cluster operated by the Laboratory Computing Resource Center at Argonne National Laboratory, and the assistance of Dr. Sean Finch at TUNL in the analysis of targets used in this research.
Appendix A Activation
The reaction rate (R) is given by,
where is the cross section, is the number of atoms in the target, and is the beam flux of incident particles.
An example of a radioisotope production curve is provided in Figure 13. The isotope is being produced faster than it can decay during the time from 0 to when the target is being irradiated.
A.0.1 197Au(,n)196Au reaction
The 196Au nucleus has three different long-lived states; e.g. the ground, and isomeric states denoted by g (6.17 d), (8.1 s) and (9.6 h). The two isomeric levels decay 100 to the ground state. The number of nuclei produced during the irradiation time can be calculated using the formulae below,
where the R’s and ’s are the reaction rates and decay constants, respectively. is the number of nuclei in the th state. The solutions for the number of nuclei in the and isomeric states are obtained from the following equations.
The number can be approximately constant, since due to the extremely short half-life compared to that of ground and isomeric states. This means . In other words, we can determine the cross sections for the production of and , separately. Using the above solutions and equation 12, we find the solution for ,
where . When the irradiation is finished, the nuclei will decay with no production term. The number of nuclei decreases with the respective half-lives,
with,
The total number of events with measurement time() will be,
Using these results and the relation between the number of events and the cross section in equation 11, the cross section for producing 196Au in the 2nd isomer state is,
Similarly, the combined cross section for ground and isomeric state can be obtained,
with,
A.0.2 reaction
The cross section for producing is determined in the same way as Equation 14,
Appendix B Beam flux simulation results
Beam flux simulation results for each segment and target are listed in the Table 8.
| E | 14 MeV | 16 MeV | 18 MeV | 22 MeV | |||||||||
| seg | tgt. | flux () | flux () | flux () | flux () | ||||||||
| 1 | Au | 13.956 | 0.181 | 23.960 | 15.944 | 0.227 | 24.400 | 17.907 | 0.243 | 24.920 | 21.828 | 0.483 | 24.680 |
| TiO2 | 13.956 | 0.182 | 23.080 | 15.944 | 0.226 | 23.430 | 17.906 | 0.243 | 23.840 | 21.825 | 0.497 | 23.536 | |
| Zn | 13.956 | 0.181 | 22.960 | 15.944 | 0.226 | 23.300 | 17.906 | 0.243 | 23.710 | 21.824 | 0.500 | 23.404 | |
| Os | 13.956 | 0.182 | 22.670 | 15.944 | 0.227 | 23.090 | 17.906 | 0.243 | 23.520 | 21.823 | 0.504 | 23.204 | |
| Au | 13.956 | 0.182 | 22.300 | 15.944 | 0.227 | 22.720 | 17.906 | 0.243 | 23.050 | 21.822 | 0.510 | 22.709 | |
| 2 | Au | 13.477 | 0.303 | 23.320 | 15.303 | 0.430 | 23.180 | 17.100 | 0.486 | 23.160 | 20.903 | 0.692 | 23.285 |
| TiO2 | 13.477 | 0.303 | 22.280 | 15.303 | 0.430 | 21.990 | 17.100 | 0.486 | 22.300 | 20.901 | 0.697 | 22.363 | |
| Zn | 13.477 | 0.303 | 22.180 | 15.302 | 0.430 | 21.890 | 17.100 | 0.486 | 22.220 | 20.901 | 0.697 | 22.276 | |
| Os | 13.477 | 0.303 | 21.960 | 15.302 | 0.430 | 21.730 | 17.100 | 0.486 | 22.000 | 20.900 | 0.699 | 22.053 | |
| Au | 13.477 | 0.303 | 21.650 | 15.302 | 0.430 | 21.440 | 17.100 | 0.486 | 21.730 | 20.900 | 0.700 | 21.763 | |
| 3 | Au | 12.915 | 0.414 | 19.270 | 14.581 | 0.468 | 18.830 | 16.184 | 0.637 | 18.610 | 19.854 | 0.824 | 18.871 |
| TiO2 | 12.914 | 0.414 | 18.230 | 14.581 | 0.468 | 17.880 | 16.184 | 0.637 | 17.720 | 19.852 | 0.827 | 17.903 | |
| Zn | 12.914 | 0.414 | 18.140 | 14.581 | 0.468 | 17.780 | 16.184 | 0.637 | 17.620 | 19.852 | 0.828 | 17.807 | |
| Os | 12.914 | 0.414 | 17.980 | 14.581 | 0.468 | 17.630 | 16.183 | 0.637 | 17.430 | 19.851 | 0.828 | 17.609 | |
| Au | 12.914 | 0.415 | 17.410 | 14.580 | 0.468 | 17.100 | 16.183 | 0.638 | 17.090 | 19.851 | 0.829 | 17.235 | |
| E | 24 MeV | 26 MeV | 31 MeV | ||||||||||
| seg | tgt. | flux () | flux () | flux () | |||||||||
| 1 | Au | 24.115 | 0.573 | 24.619 | 26.232 | 0.655 | 25.179 | 31.053 | 0.843 | 26.180 | |||
| TiO2 | 24.111 | 0.596 | 23.525 | 26.227 | 0.688 | 24.089 | 31.040 | 0.933 | 24.850 | ||||
| Zn | 24.110 | 0.602 | 23.389 | 26.225 | 0.697 | 23.958 | 31.037 | 0.948 | 24.710 | ||||
| Os | 24.109 | 0.609 | 23.159 | 26.222 | 0.714 | 23.620 | 31.034 | 0.968 | 24.491 | ||||
| Au | 24.107 | 0.619 | 22.741 | 26.219 | 0.732 | 23.168 | 31.007 | 1.125 | 23.068 | ||||
| 2 | Au | 22.987 | 0.830 | 23.308 | 24.912 | 0.948 | 23.239 | 29.245 | 1.264 | 23.080 | |||
| TiO2 | 22.982 | 0.844 | 21.974 | 24.906 | 0.969 | 22.019 | 29.237 | 1.295 | 22.083 | ||||
| Zn | 22.982 | 0.845 | 21.870 | 24.906 | 0.970 | 21.909 | 29.237 | 1.297 | 21.994 | ||||
| Os | 22.981 | 0.846 | 21.691 | 24.905 | 0.973 | 21.665 | 29.234 | 1.305 | 21.763 | ||||
| Au | 22.980 | 0.847 | 21.369 | 24.904 | 0.977 | 21.316 | 29.220 | 1.362 | 20.687 | ||||
| 3 | Au | 21.718 | 0.986 | 18.757 | 23.431 | 1.131 | 18.404 | 27.239 | 1.509 | 17.814 | |||
| TiO2 | 21.714 | 0.993 | 17.694 | 23.425 | 1.144 | 17.238 | 27.230 | 1.532 | 16.818 | ||||
| Zn | 21.714 | 0.993 | 17.599 | 23.425 | 1.144 | 17.150 | 27.229 | 1.533 | 16.726 | ||||
| Os | 21.714 | 0.994 | 17.433 | 23.424 | 1.145 | 16.980 | 27.228 | 1.535 | 16.530 | ||||
| Au | 21.712 | 0.997 | 16.857 | 23.421 | 1.151 | 16.347 | 27.217 | 1.563 | 15.582 |